Circular No. 7894

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002cj, 2002ck, 2002cp, 2002cs T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cj (cf. IAUC 7882), obtained by M. Calkins on May 2.43 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ic supernova, similar to SN 2002cg (cf. IAUC 7881). R. Chornock, W. D. Li, and A. V. Filippenko, University of California, Berkeley, write that inspection of CCD spectra (range 330–1000 nm), taken on May 7 with the Shane 3-m reflector at Lick Observatory, also reveals that SN 2002cj is a type-Ic supernova.

Chornock et al. note that SN 2002ck (IAUC 7884) is of type Ia, with a spectral-feature age of 5 ± 2 days past maximum light on May 7.

Matheson et al. further report that a spectrum of SN 2002cp (cf. IAUC 7887), obtained as above by P. Berlind on May 3.16 UT, shows it to be a type-Ib or type-Ic supernova. The spectrum consists of a blue continuum with several absorption features, including calcium and what may be incipient helium lines. Chornock et al. agree that 2002cp (IAUC 7887) is a young supernova of type Ib or Ic. Further monitoring is encouraged to determine whether or not the He I absorption lines that are characteristic of type-Ib supernovae develop.

M. Riello, S. Benetti, G. Altavilla, A. Pastorello, and M. Turatto, Osservatorio Astronomico di Padova; E. Cappellaro, Osservatorio Astronomico di Capodimonte; and P. Ruiz-Lapuente, Barcelona University, communicate that a spectrum (range 325–800 nm) of SN 2002cs (cf. IAUC 7891), obtained on May 7 with the Telescopio Nazionale Galileo (+ Dolores spectrograph), shows it to be a type-Ia supernova 2 ± 2 days before maximum light (from spectral features); assuming the NASA/IPAC Extragalactic Database (NED) recession velocity of 4728 km/s for the host galaxy, the supernova expansion velocity is ∼ 15700 km/s from the Si II 635.5-nm absorption. Matheson et al. confirm that a spectrum of SN 2002cs, obtained by Calkins on May 7.38, shows it to be a type-Ia supernova near maximum; also adopting the NED recession velocity for NGC 6702, they find the supernova expansion velocity to be ∼ 15900 km/s for Si II (rest 635.5 nm), and the spectral-feature age of the supernova was determined to be 1 ± 2 days before maximum light, for which the high expansion velocity is unusual. Chornock et al. remark that inspection of CCD spectra, taken as above at Lick, reveals that SN 2002cs is a supernova of type Ia, closely resembling SN 1994D near maximum light.

2002 May 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green